If a computer has 64-bit addresses and each address corresponds to one byte, how many gigabytes of memory can the computer address? (Show the math)

The past answers for this same question are either incorrect or do not show the math.
Answer #1

******************************************************************************************
Please Upvote the answer as it matters to me a lot :)
*****************************************************************************************
As per HomeworkLib expert answering guidelines,Experts are supposed to answer only certain number of questions/sub-parts in a post.Please raise the remaining as a new question as per HomeworkLib guidelines.
******************************************************************************************

let's suppose you have 3 bits how many numbers can you represent using these 3 bits?

1st bit 0 or 1 = 2 choices

2nd bit 0 or 1 = 2 choices

3rd bit 0 or 1 = 2 choices

so total 2*2*2 = 8 numbers

in the same way, if you have n bits then you can represent 2*2*2*.....2(n times )= 2n numbers

given each byte is represented by one 64 bit address

total bytes = total number that 64 bits can represent = 264

we know 210 bytes = 1KB

220= 1 MB

230= 1GB

total memory the computer address is 264/ 230 GB = 234 GB